ISBN-0-07-873823-7

one number is 12 less than three times another number. their sum is 188. write two equations that would be used to solve the system.

Answer by checkley71(8403)

You don't need 2 equations for this one. 1 equation will do.
letx be 1 number & (3x-12) for the other number.
x+(3x-12)=188
4x=188+12
4x=200
x=200/4
x=50 answer.
proof
50+3*50-12=188
50+150-12=188
200-12=188
188=188